7. (a) A car needs a battery with the following specifications (12V and 10 Amps). A...

7.

(a) A car needs a battery with the following specifications (12V and 10 Amps). A flashlight battery has the following specifications (1.5 V & 0.5 Amp) . If you plan to replace a car battery with flash light batteries What is the minimum number of batteries required? How many of them will be in series and how many in parallel ?--------         (15)

(b)    Can a charge particle move through a magnetic field without experiencing any force? If your answer is “yes” then explain how it would be possible.

      If your answer is “no” then explain why it won’t be possible. -------    (5)

(c) During a power outage why it is recommended that you switch your heavy duty devices OFF and wait for the power to be back before turning them ON ----------------             (10)
Answer #1

car battery voltage= 12 volts

cuurent= 10 A

flash light battery voltage= 1.5 v

so number of batteries required in a row in series= 12/1.5 =8

now number of rows in parallel =10/0.5 = 20

so total number of batteries= 20 x8 = 160 batteries

(b)yes it is possible

magnetic force= Fm = q v b sin theta

if theta= 0o or 180o then Fm= 0 as sin0 = 0 and sin 180 =0

so if a charged body is moving parallel to magnetic filed, it will not experience any force.

(c) it is recommend to switch off the heavy duty devices during outage to protect them from huge and sudden currents when the power comes back.